OpenStudy (anonymous):

ive got 17 questions

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

x/5-x/6=1/3

OpenStudy (anonymous):

x = 10

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

c-1/x+d=cd

OpenStudy (anonymous):

x = 1/(cd - c - d)

OpenStudy (anonymous):

do you want me to explain these at all?

OpenStudy (anonymous):

that is wrong

OpenStudy (anonymous):

Solve for x.

OpenStudy (anonymous):

c-1 is over x